OpenStudy (anonymous):

Identify the relative maximum value for the function shown below.? g(x)=7/x^2+5

OpenStudy (baldymcgee6):

is it g(x) = 7/(x^2 +5) ?

OpenStudy (anonymous):

yes

OpenStudy (baldymcgee6):

okay, think of it this way... what can you put in for x to give you the biggest output?

OpenStudy (baldymcgee6):

if you put in a really big x value in, what will you get? 7/(really big number) = small number

OpenStudy (baldymcgee6):

so the way to get the biggest output is to put in a really small number in for x

OpenStudy (anonymous):

like 0?

OpenStudy (baldymcgee6):

exatly zero. Because when you put 0 in for x, you 7/5 which is the biggest possible number, also check out the graph at x =0 http://screencast.com/t/A1yoHw8EM0K0

OpenStudy (anonymous):

differentiate use maximum test

OpenStudy (anonymous):

how? i'm so confused right now

OpenStudy (baldymcgee6):

@classof2013 is this for pre-calculus?

OpenStudy (anonymous):

yes it is

OpenStudy (baldymcgee6):

okay, then yes, we can use the approach suggested by gomathi

OpenStudy (baldymcgee6):

Lets differentiate g(x)=7/x^2+5

OpenStudy (anonymous):

how do you do that

OpenStudy (anonymous):

wait wouldn't it be 14 then?? or 0

OpenStudy (baldymcgee6):

Your're getting ahead of yourself, first lets differentiate g(x)=7/x^2+5 so this is a quotient, so lets use the quotient rule. Do you know what that is and how to use it?

OpenStudy (anonymous):

no i don't

OpenStudy (baldymcgee6):

Have you been taught differentiating rules yet?

OpenStudy (anonymous):

well it's all on the laptop and they don't really teach us much

OpenStudy (baldymcgee6):

Okay, then, lets learn it! when we have a function in the form of g(x) = f(x)/h(x), ie, a quotient, we will apply the quotient rule.

OpenStudy (baldymcgee6):

the quotient rule will give us g'(x) = (f'h-fh')/h^2

OpenStudy (anonymous):

then after that?

OpenStudy (baldymcgee6):

then we get our derivative and equate it to zero, ie. our slope=0 (a horizontal line) http://screencast.com/t/NKOvnuGEH

OpenStudy (anonymous):

how do we know which one is maximum & minimum?

OpenStudy (baldymcgee6):

look at the graph

OpenStudy (baldymcgee6):

if you look at the link I gave you, you will notice the maximum right away.

OpenStudy (anonymous):

yeah isnt it the line right at top?

OpenStudy (baldymcgee6):

that horizontal line is the slope of zero

OpenStudy (anonymous):

i get it kind of

OpenStudy (baldymcgee6):

good, let me know if there is anything else..